OpenStudy (anonymous):

Determine which ordered pair is a solution of y = -5x - 1 . a. (9, –2) b. (5, 24) c. (–1, 6) d. (–4, 19)

OpenStudy (anonymous):

so which part are you stuck on. did you try plugging in the x for each of the coordinate pairs and then see if both sides are equal like for example the first coordinate pair if you plug in 9 for x and -2 for y, you should get -2 = -5(9) -1 ; which equals -2=-45 -1; that equals -2 = -46 which is not true right so we can cross that answer off. let me know if you still need help